OpenStudy (maddie82502):

Which ratio is equivalent to 4 : 13? A. 8 : 17 B. 16 : 52 C. 12 : 33 D. 20 : 55

OpenStudy (maddie82502):

@americansniper12 @Brooklyn_Boiii

OpenStudy (anonymous):

compare the options given with \[4:13=\frac{ 4 }{ 13 }\]

OpenStudy (anonymous):

\[\frac{ 4 }{ 13 }\neq \frac{ 8 }{ 17 }\]

OpenStudy (anonymous):

Therefore,option A is wrong

OpenStudy (dakid88):

the answer is B

OpenStudy (lord2clash):

the answer is B. because 4 goes into 16 four times and 13 goes into 52 four times.

OpenStudy (maddie82502):

thanks

OpenStudy (lord2clash):

the others either 1 number only goes in or neither do.

OpenStudy (anonymous):

Option B\[\frac{ 4 }{ 13 }=\frac{ 16 }{ 52 }=\frac{ 4 }{ 13 }\]Option C\[\frac{ 4 }{ 13 }\neq\frac{ 12 }{ 33 }\]Option D\[\frac{ 4 }{ 13 }\neq \frac{ 20 }{ 55 }\neq \frac{ 4 }{ 11 }\]

OpenStudy (anonymous):

yep it is B

OpenStudy (dakid88):

hey i said the answer how come i didnt get a medal

OpenStudy (lord2clash):

@dakid88 you have to explain your answers aswell as giving them (altho it is better to make the person get the answer themselves by giving them the information all the way until the answer have them get the answer)